OpenStudy (anonymous):

What is the sum? 7/3+(-3/8) ((The subtraction symbol is supposed to be for the entirety of 3/8, not just 3, to clarify. The two answers below, C and D, are negative fractions as well)) A. 65/24 B. 47/24 C. -4/5 D. -5/4

OpenStudy (anonymous):

\[\large \frac{7}{3}~+~(-\frac{3}{8})~?\]

OpenStudy (anonymous):

That is what it looks like right?

OpenStudy (anonymous):

Yup, that's the one. How on earth do you- oh is that on the "equation" tab at the bottom? Ahhaha... I'm slow. Yes that's what I mean.

OpenStudy (anonymous):

Nah, I wrote it out using latex.

OpenStudy (anonymous):

what's what?

OpenStudy (anonymous):

But anyways, \[\large +(-x)~=~-x\] so \[\large +(-\frac{3}{8})~=~-\frac{3}{8}\] and btw, if you want to learn latex check out http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A)

OpenStudy (anonymous):

I.. Still don't quite understand this. I will learn this latex business though.

OpenStudy (anonymous):

What I'm trying to say is \[\large (\frac{7}{3}~+~(-\frac{3}{8}))~=~(\frac{7}{3}~-~\frac{3}{8})\]

OpenStudy (anonymous):

So the next step is to get the a common denominator. Know how to do that?

OpenStudy (anonymous):

Oh yeah, I can do that. But with this problem it's kinda difficult for me and I don't know how to apply it.

OpenStudy (anonymous):

nevermind I got it it's 47/24